### v3

# A Comprehensive Survey on Data Augmentation

Abstract
Data augmentation is a series of techniques that generate high-quality artificial data by manipulating existing data samples. By leveraging data augmentation techniques, AI models can achieve significantly improved applicability in tasks involving scarce or imbalanced datasets, thereby substantially enhancing AI models’ generalization capabilities. Existing literature surveys only focus on a certain type of specific modality data, and categorize these methods from modality-specific and operation-centric perspectives, which lacks a consistent summary of data augmentation methods across multiple modalities and limits the comprehension of how existing data samples serve the data augmentation process. To bridge this gap, this survey proposes a more enlightening taxonomy that encompasses data augmentation techniques for different common data modalities by investigating how to take advantage of the intrinsic relationship between and within instances. Additionally, it categorizes data augmentation methods across five data modalities through a unified inductive approach.
Index Terms:
Data augmentation, Data-centric taxonomy, Multi-modality
I Introduction
With the rapid development of Artificial Intelligence (AI) in the past decades, AI methods have shown their superiority over human beings and other traditional approaches across most tasks, from universal tasks in our daily lives such as image recognition [1] and text translation [2] to complicated scientific research tasks such as drug-drug interaction prediction [3] and single-cell RNA sequencing analysis [4]. Recently, powered by Latent Diffusion Models [5] and Transformer [6], Stable Diffusion [5] and ChatGPT [7] have greatly changed human beings’ working, living, and entertainment. The success of AI products is usually attributed to AI models’ in-depth understanding of the accumulated data, thereby intrinsically uncovering data patterns and learning data-task correlations. The performances of AI models will be affected by the quantity and quality of training data, e.g. models trained with limited data suffer from over-fitting with significantly degenerative performance on testing datasets, and models trained with imbalanced samples will lead to poor generalization ability. In most situations, researchers have to overcome difficulties caused by data scarcity and distribution imbalance. An intuitive solution to these problems is to acquire more data, but data are limited or difficult to collect in many cases, and labeling the data is yet another labor-intensive work.
To solve these problems, data augmentation has been extensively applied and proven to be effective and efficient [8]. The core idea of data augmentation is to artificially enlarge the training dataset by creating modified copies of existing data. It also introduces more diversity and fills the gap between training datasets and real-world applications. As technology continues to evolve, a variety of techniques are applied to augment data samples. Some simple approaches only randomly mask part of the data [9, 10, 11], while more sophisticated augmentation schemes involve generative adversarial networks [12] or reinforcement learning agents [13].
An ideal survey on data augmentation should be modality-independent, because only in this way can it focus on the intrinsic mechanisms of data augmentation, regardless of the data type, and thus provide insight into the nature of data augmentation. Existing literature surveys summarize data augmentation methods from different perspectives [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. As summarized in Table. I, each of them concentrates on a certain data modality. Most of them categorize data augmentation methods with modality-specific (e.g. image modality and text modality) or operation-centric (e.g. feature scaling operation, data perturbation operation) taxonomies. These surveys reflect the development of data augmentation and its application in different learning scenarios. However, they fail to cover data augmentation methods across data types, uncover their common patterns despite different modalities, and thus limit readers’ understanding of the essence of data augmentation. To fill this gap, we investigate data augmentation methods comprehensively, across the five most popular data modalities. We analyze how data augmentation methods utilize different sample numbers and how they leverage different components of the data information. Then, we propose a unified taxonomy that focuses on the way data augmentation leverages information from samples. Finally, we investigate up-to-date literature and categorize them by our proposed taxonomy.
The main contributions of this survey are as follows:
  • We propose a novel modality-independent taxonomy from a data-centric perspective that accommodates data augmentation techniques for all modalities consistently and inductively;
  • To the best of our knowledge, this is the first survey that covers data augmentation techniques across five data modalities, i.e. image, text, graph, tabular, and time-series modalities;
  • We investigate how information is consistently contained in each of these data modalities and can be utilized for data augmentation;
  • We include and categorize up-to-date literature in data augmentation.
TABLE I: A comparison of data modalities and taxonomies discussed by existing surveys and ours.
Data Coverage Taxonomies
 Image Text Graph Tabular Time-series  Modality-specific Modality-independent Operation-centric Data-centric
IIPhDW 2018 [14]
J Big Data 2019 [15]
IJCNLP 2021 [18]
J Big Data 2021 [19]
IJCAI 2021 [27]
Plos one 2021 [29]
Artif Intell Rev 2022 [16]
Array 2022 [17]
Comput Surv 2022 [20]
KDD 2022 [21]
arXiv 2022 [22]
arXiv 2022 [23]
TNNLS 2022 [24]
J Big Data 2022 [25]
J Big Data 2023 [26]
Neural Comput Appl 2023 [28]
arXiv 2024 [30]
arXiv 2024 [31]
arXiv 2025 [32]
Ours
II Background
II-A Early Evolution
Despite the fact that data augmentation only became popular in no more than a decade, some of its embryonic forms were proposed quite early. The concept of data augmentation was conceived at the very beginning of Deep Learning. For example, the use of random distortion can be found in LeNet [33], where ninefold the distorted images are added to the dataset to verify that increasing the size of the training set can effectively reduce test error. Since then, data augmentation has gradually been recognized as a best practice for training Convolutional Neural Networks (CNNs) [34, 35]. AlexNet [36] explicitly employs several data augmentation techniques to reduce overfitting. It augments the datasets by extracting patches from images and altering the color intensity. It also applies dropout, which randomly sets the output of some neurons to zero to prevent their co-adaption. Although dropout is often regarded as an approach to regularization, it has a similar idea to data augmentation and indeed inspires a set of augmentation methods [9]. These applications of data augmentation mainly focus on increasing the size of the training dataset and introducing diversity to it. SMOTE [37] is another early and representative work on data augmentation but from a different perspective. Focusing on addressing the class imbalance problem, it suggests that oversampling the minority class can achieve better classification performance when categories are not equally represented in the dataset.
II-B Existing Surveys
Given the rapid development of CNNs and their wide application in image-related tasks, most early data augmentation methods are proposed for image data. Also, it is natural to augment image data because many data augmentation methods suit the traits of CNNs very well, such as translation invariance. Published in 2018, [14] is one of the first surveys on data augmentation. It briefly discusses the effects of traditional and neural-based augmentation methods. In 2019, data augmentation usage on image data is thoroughly studied in [15]. It categorizes image data augmentation methods into basic image manipulation and deep learning approaches, with a wide coverage. It also introduces how meta-learning selects and combines basic data augmentation operations. [16] is a recent survey on image data augmentation published in 2022. It formalizes commonly used operations for image data augmentation with equations and illustrates their effects with examples. In the same year, [17] also elaborates on image data augmentation with plenty of examples. It adopts taxonomies similar to [15], but includes more modern approaches and evaluates the improvements in model performance.
Data augmentation on text data has not been thoroughly researched as early as image data, possibly due to the discrete and correlated nature of text components. Still, several augmentation methods are proposed for or transplanted to text data. [18] is among the first surveys on text data augmentation. It adopts a simple taxonomy but gives more emphasis on the applications. Built on previous work on image data augmentation, [19] continues to investigate how data augmentation is used to enrich text datasets, and methods are classified as symbolic augmentation and neural augmentation. [20] categorizes text data augmentation methods by identifying whether they are used on the data space or feature space. It also presents the performance improvements of different methods. [31] focuses on the application of data augmentation for training Large Language Models. [32] discusses how Multimodal Large Language Models can be used for augmenting text, image, and speech data.
Following the success of data augmentation methods for image and text data, there is a growing trend in augmenting graph data, too. Since the graph represents data in a relatively complex form compared with image and text, there are more opportunities for data augmentation. This also leads to more possible ways to categorize graph data augmentation methods. For example, [21] views graph data augmentation from the employed techniques and application scenarios, while [22] discusses graph data augmentation in terms of its data modality, task level, and whether the augmentation operation is rule-based or learned. [23] tries to elaborate on all feasible taxonomies for graph data augmentation and uses detailed schematic illustrations to explain the overview framework as well as the idea of some typical methods.
Tabular data is common in reality, but there is not as much room for data augmentation as other data types because tabular data lacks some features for augmentation to utilize. Some early surveys only discuss augmentation techniques in the context of deep learning on tabular data [24], and others mainly focus on how to generate tabular data [25, 26]. Recently, [30] presented a dedicated and thorough review on tabular data augmentation, focusing on how generative models and retrieval techniques contribute to tabular data augmentation and discussing their future trends.
With its growing application in scenarios such as speech recognition and IoT, time-series data is attracting more attention. [27] discusses basic and advanced augmentation methods. [28] focuses on generative approaches. [29] gives more discussion on mixing time-series data.
Till now, data augmentation techniques for different data types have been exhaustively discussed; however, there has not been a comprehensive survey that summarizes data augmentation for all types of data. Such a survey can be profitable because data augmentation techniques for different modalities share some common methodologies when leveraging information from existing data, such as altering some features of a sample or mixing values between multiple samples. Summarizing these similarities can help reveal the common pattern of data augmentation. A comprehensive survey does not mean simply combining existing surveys because the proposed taxonomies in these surveys view data augmentation from different angles and are incompatible with each other. Some taxonomies contrast basic operations to advanced approaches; some categorize methods with modality-specific terms. Besides, these taxonomies provide limited insight into what happens to the information. Indeed, huge gaps lie between operations for data modalities due to their distinctive nature; these operation-based or modality-specific taxonomies reflect the development of data augmentation techniques. However, from such taxonomies, we cannot grasp the essence of data augmentation and lack perception of the big picture of data augmentation. Hence there is a need for a comprehensive survey that covers data augmentation techniques for all data modalities with an all-embracing and data-centric taxonomy.
III Taxonomy
Before presenting the taxonomy, we formalize data augmentation as follows. Given a labeled dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$, where $\mathbf{X}$ stands for the data, and $\mathbf{y}$ stands for the labels, data augmentation can be represented by a certain function $f_{\theta}$, such that the augmented dataset $\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ is derived by:
$f_{\theta}:\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}\rightarrow\mathcal{\tilde{D}}_{L}=\{\mathbf{\tilde{X}},\mathbf{\tilde{y}}\}$ (1)
This representation also applies to an unlabeled dataset $\mathcal{D}_{U}=\{\mathbf{X}\}$ or a partially labeled dataset $\mathcal{D}_{P}=\{\mathbf{X}_{L}\cup\mathbf{X}_{U},\mathbf{y}_{L}\}$, but for simplicity, we only use a labeled dataset to explain the taxonomy. Some data augmentation approaches are used in the input data space, while others are used in the latent feature space. This section uses $\mathbf{X}$ as a space-agnostic representation.
In this survey, we try to propose a two-tier taxonomy from a data-centric perspective that can be applied to all data modalities. The main consideration of our taxonomy is whence the information in the augmented data comes. It differentiates data augmentation techniques by asking two research questions: RQ1: How many samples are used to generate each new sample? RQ2: Which part of the information is used to generate new data? The answers to these questions constitute the taxonomy hierarchy which we present below.
By answering RQ1, we can divide data augmentation approaches into Single-instance Level Augmentation, Multi-instance Level Augmentation, and Dataset Level Augmentation, which build the first tier of the taxonomy in our survey.
Figure 1: A taxonomy for data augmentation (left) and how information in each data modality is utilized (right).
Single-instance Level Augmentation. This type of data augmentation method transforms one data sample without referring to others to create new data. It introduces a certain perturbation to the data, such as masking and noise addition. The information of new data comes from exactly one sample in the original dataset, and each new sample often lies around its original sample in the feature space. We formalize single-instance level augmentation as follows, where $\epsilon(\mathbf{x}_{i})$ stands for a certain perturbation related to the original data $\mathbf{x}_{i}$:
$\mathbf{\tilde{x}}=\mathbf{x}_{i}+\epsilon(\mathbf{x}_{i}),\qquad\tilde{y}=y_{i}$ (2)
We further answer RQ2 to build the second tier of the taxonomy. To answer this question, we must consider how the information is contained in the data. Typically, data is made up of a certain type of element that carries unit information with a value, such as the coloring of a pixel or the wording in a sentence. For most data, there is also some structural information that connects these elements in a certain way, such as the positional relationship of pixels or the syntax of a sentence. The key to designing data augmentation methods is to think of how to wisely leverage the two types of information. So, we can categorize single-instance level augmentation according to which part of the information is used for data augmentation:
  • Value-based Transformation, which perturbs the value that an element carries.
  • Structure-based Transformation, which perturbs elements’ structural relationships.
Multi-instance Level Augmentation. This type of data augmentation method makes use of multiple data to acquire new ones. Information from different data samples is combined by operations such as interpolation and concatenation. New samples derived from these methods often lie between the source data in the feature space. We formalize multi-instance level augmentation as follows, where $i$ and $j$ are two data samples from the original dataset, and $\lambda$ is a scalar factor that often lies between $[0,1]$:
$\mathbf{\tilde{x}}=\lambda\cdot\mathbf{x}_{i}+(1-\lambda)\cdot\mathbf{x}_{j},\qquad\tilde{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j}$ (3)
Now we answer RQ2 for multi-instance level augmentation. Similarly to single-instance level augmentation, this type of data augmentation is also conducted utilizing either value or structure information, and the detailed classification is:
  • Value-based Mixture, which arithmetically mixes the values of multiple data.
  • Structure-based Combination, which pieces multiple data (or parts of them) together.
Dataset Level Augmentation. This type of data augmentation method does not explicitly make use of one or multiple data but rather uses the entire dataset. It comprehensively learns the features and abstractions of the dataset, and samples completely new data from a learned distribution of the dataset. The generated data should also fall into this distribution. We formalize dataset level augmentation as follows, where $P(\cdot)$ stands for a probability distribution learned from the original dataset $\mathcal{D}_{L}=\{\mathbf{X},\mathbf{y}\}$:
$\vspace{-1.5mm}\mathbf{\tilde{x}}\sim P(\mathbf{X}),\qquad\tilde{y}\sim P(\mathbf{y}|\mathbf{\tilde{x}})$ (4)
The answer to RQ2 for dataset level augmentation is different from the two categories mentioned above. When generating new data based on dataset distributions, most methods simultaneously consider value and structural information. Some methods also introduce external knowledge when learning from the source dataset. So, we have the following classification:
  • Vanilla Generation, where the generation process only resorts to the existing dataset.
  • Exogenous Generation, which uses external resources like other datasets or expert knowledge.
We have presented a general taxonomy for data augmentation at a high level with abstract and conceptual formalization. We organize this survey in accordance with this taxonomy, from Section IV to Section VI. We elaborate on how the proposed taxonomy applies to each data modality and review some typical data augmentation methods.
IV Single-instance Level Augmentation
Single-instance level augmentation methods perturb one data sample’s value or structure information to obtain the augmented data. Different data modalities represent value information with different elemental components and organize these components with different types of structure.
For image data, the elemental components are pixels. A pixel carries unit information that uses a numerical value to represent the color. Each pixel also has spatial relationships with other pixels, so, images are well-structured. Single-instance level augmentation for image data revolves around the perturbation of the image pixels’ colors and spatial relationships. For text data, the elemental components of text data are words (or sometimes, tokens). A word often expresses minimal meanings. Words are organized into sentences following syntactic rules. The sentence is a sequential representation since one word comes after another, but there is also a tree-structured grammatical dependency behind the sentence. Single-instance level augmentation for text data often changes the wording or reshapes the syntax of a sentence. A graph is represented by $\mathbf{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ stands for the vertex set, and $\mathcal{E}$ stands for the edge set. The topology structure can be alternatively represented by an adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, where $\mathbf{A}_{ij}$ is $1$ if there is an edge connecting vertex $i$ and vertex $j$, and is 00 otherwise. In a deep graph learning scenario, the graph also contains a feature matrix $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ that represents the attributes of the nodes. Hence another representation for a graph is $\mathbf{G}=\{\mathbf{A},\mathbf{X}\}$. Single-instance level augmentation for graph data is to perturb the values in the feature matrix $\mathbf{X}$ of nodes or the topology structure represented by the adjacency matrix $\mathbf{A}$. A table is made up of cells, which store numerical or categorical values that represent attributes of the described entities. Cells are organized into columns and rows, but this structural information is trivial, and switching the sequence of the columns and rows does not affect the information conveyed by the table. So, most data augmentation methods leverage the value information. Still, it is also possible to perturb the structure information of a table. Time-series data represent sequential values within a period, and each timestamp is considered an elemental component. Data augmentation methods for time-series data either perturb the value at each timestamp or perturb the sequence of timestamps.
That is, single-instance level augmentation methods focus on utilizing the values and structures of these components.
IV-A Value-based Transformation
IV-A1 Color Perturbation
Pixel Erasing masks out all color information in an area of the data; Photometric Transformation imposes a soft adjustment to change the appearance of the image, but the essential information remains visible.
Pixel Erasing
As illustrated in Fig. 2b, the idea of pixel erasing is to mask some information from the original input space or the feature space. Cutout [9] and random erasing [10] first selects a rectangle region and erases all pixels in that region. The erased region is then filled with a color or Gaussian noise. Random Erasing typically adds one or two big rectangular masks to the image, and some other methods deploy more flexible implementations. Hide-And-Seek [38] randomly removes multiple parts of the image so that the model has to learn from multiple parts of the image other than the most discriminating part. Similarly, the Grid-Mask method [39] generates a set of squared masks in the grid layout. It can prevent excessive deletion and reservation and achieve a reasonable balance between deleted and reserved regional information. In most research, random erasing is used in the spatial domain, while REF [40] extends the use of random erasing to the frequency domain. It uses the discrete Fourier transform to process the image, applies random erasing on the image’s frequency domain, and uses the inverse discrete Fourier transform to recover the image to the spatial domain.
Photometric Transformation
Augmentation that changes the overall color and style of the image is called photometric transformation, as shown in Fig 2c. The simplest way is to adjust the brightness of the image. Color inversion replaces the pixels with their “opposite” color. Color casting applies different operations in different channels, for example, increasing the image’s red channel while keeping the green channel and blue channel untouched. Noise injection adds Gaussian or other types of noises to the images. Greyscaling transforms a colored picture into black-and-white by combining channels for different colors, forcing the model to focus on the shapes or edges of target objects instead of their colors. Histogram equalization [41, 42] enhances the image by enhancing its contrast, and white balancing simulates alternative light cast on the object. Kernel filters are matrices that slide across and are convoluted with the original image to generate new images, either sharpening or blurring the image.
Figure 2: A conceptual demonstration of (a) Original Image, (b) Pixel Erasing, (c) Photometric Transformation, (d) Image Cropping, (e) Geometric Transformation, (f) Policy-based Data Augmentation (g) Prompt-based Image Editing.
IV-A2 Wording Perturbation
Text data augmentation of this type uses word-level manipulation. It applies changes to the choice of words or adds or removes words. Some examples are presented in Table II.
Token Replacement
Token replacement refers to using alternative wording in a given sentence. EDA [11] proposes a set of text augmentations, including synonym replacement. It randomly selects several words from a sentence, substituting them with their synonyms, i.e. words that mean exactly or nearly the same. [43] uses WordNet [44], a lexical database to find synonyms for verbs and nouns in sentences and thus make replacements. It also derives word embeddings by Word2Vec [45] and replaces a word with others that have high cosine similarity. Another possibility is to use a knowledge graph. A knowledge graph describes the relationships of entities, and by utilizing the “is equivalent” relation, we can find a series of synonyms for a given word.
Token Addition
Token addition refers to adding tokens to the original text, either introducing junk tokens as noise or enriching the sentences and making them more informative. The random insertion, proposed in EDA [11], inserts a random synonym of a random word at a random position in the sentence. The words added in the EDA play the role of data noising and may carry either little or harmful information. As a variant of EDA, the technique of AEDA [46] only inserts punctuation marks.
Token Deletion
Random deletion [11] randomly removes some words from a sentence. In [47], noising techniques are used to avoid overfitting, and the randomly selected word is replaced by a blank placeholder token “_”, rather than simply being removed. In [48], apart from word dropout which is similar to random deletion, it also uses word blanking to replace a word with a special “¡BLANK¿” placeholder.
IV-A3 Node Attributes Perturbation
Graph nodes often contain rich attributes in the feature matrix. This matrix is generally similar to tabular data, so augmenting node attributes does not need to consider the specialty of graph data.
Attribute Masking and Noise Injection
Various typical augmentation approaches on node attributes from random masking [49, 50, 51] to perturbing attributes with random noises [52] can be applied, as we do for common tabular data. These methods are unified in Eq. 5, where $\mathbf{M}\in\{0,1\}^{N\times N}$ is a binary mask, $\odot$ denotes the Hadamard product, and $\mathbf{N_{W}}\sim\mathcal{N}(0,\Sigma)$ is the Gaussian white noise.
$\mathbf{\tilde{X}}=\mathbf{X}\odot\mathbf{M}+\mathbf{N}_{W}$ (5)
Some other methods are more sophisticated. For example, [53] first learns the topology information of the graph by random walk, and uses the topology embedding together with the original features as inputs to a dual graph neural network. Similarly, LA-GNN [54] applies a local augmentation mechanism. It samples features from nodes’ neighborhoods as additional information when augmenting the attributes. A different approach is proposed in DropMessage [55]. Instead of directly modifying the attributes on the nodes, it applies random dropping in the message-passing phase.
IV-A4 Cell Value Perturbation
Tabular data is much simpler compared with other modalities, so it is hard to design complex augmentation methods that make use of spatial or sequential relationships within a data sample as we do for image or text data, and most data augmentation for tabular data perturbs the value stored in the table cells. Table Masking is a simple technique often applied for tabular data augmentation. These corrupted data are often used with the original data under a self-supervised learning framework to learn stronger representations of the dataset. Another approach to tabular data augmentation borrows ideas from Feature Engineering. Some work on Feature Engineering automates feature selection and transformation to generate optimal feature space, and the operations can also be used to augment tabular datasets.
Table Masking
VIME [56] uses a mask generator to corrupt features. In a self-supervised scenario, two pretext tasks respectively reconstruct the original data and estimate the mask vector. The model updates its weights based on the reconstruction loss and the mask vector estimation loss of the two pretext tasks. VIME also considers a semi-supervised scenario. In that case, multiple masks are generated and applied to the feature of the unlabeled samples. Despite not being labeled, the model predictions on these corrupted samples should be similar, and thus a consistency loss is calculated. For labeled samples, a supervised loss is calculated based on the prediction. The two losses jointly contribute to the back-propagation phase. SCARF [57] applies data augmentation in a contrastive learning framework. Its corruption process is similar to that of VIME but it uses InfoNCE loss to evaluate the similarity between the embeddings of the corrupted and original views and then update the model accordingly. MET [58] works in a pattern similar to VIME and SCARF, which calculates the reconstruction loss of a partially masked sample.
Feature Engineering
Feature transformation [59, 60] and feature generation techniques can effectively enlarge tabular datasets by generating tables with different feature sets [61]. Powered by reinforcement learning, GRFG [62] transforms existing features with cascading agents. It uses arithmetic operations to provide better traceability and explainability. MOAT [63] extends feature transformation to continuous space. It trains an encoder to project the feature transformation process to continuous embeddings and decodes the optimized embeddings for the best feature transformation strategies. LFG [64] and TIFG [65] leverage LLM agents to generate new features that are both explainable and effective. TFWT [66] computes weights for samples and features, and the weights are combined with the data for later tasks. Although these studies focus on feature engineering, their way of thinking and application of reinforcement learning is very enlightening for data augmentation tasks.
IV-A5 Sequence Value Perturbation
The value information of a sequence is carried on the Y-axis, represented by the amplitude at specific timestamps or frequency channels. Amplitude Perturbation augments time-series data by altering the Y-axis values, and Decomposition decomposes the value of a sequence into different parts that express the sequential values from different aspects.
Amplitude Perturbation
Jittering adds random noise to the value at each timestamp of a sequence. In [67], jittering is used to simulated additive noise $\epsilon$ of wearable sensors:
$\tilde{x}(\epsilon)=\{x_{1}+\epsilon_{1},\cdots,x_{t}+\epsilon_{t},\cdots,x_{T}+\epsilon_{T}\}$ (6)
Scaling [67] alters the data by multiplying the magnitude of the entire sequence with a scalar $\alpha$.
$\tilde{x}(\alpha)=\{\alpha x_{1},\cdots,\alpha x_{t},\cdots,\alpha x_{T}\}$ (7)
Magnitude warping [67] is similar to scaling, but it uses a smooth curve to warp the sequence so that magnitude at different timestamps is scaled with different parameters.
$\tilde{x}(\alpha)=\{\alpha_{1}x_{1},\cdots,\alpha_{t}x_{t},\cdots,\alpha_{T}x_{T}\}$ (8)
Rotation [67], or flipping, assumes the data is symmetric with respect to the X-axis, and makes the data upside down by applying a rotation matrix $R$.
$\tilde{x}(R)=\{Rx_{1},\cdots,Rx_{t},\cdots,Rx_{T}\}$ (9)
Apart from the amplitude, the phase spectrum provides another option for data augmentation. RobustTAD [68] converts the time-series data from the time domain to the frequency domain, as shown in Eq. 10, and then increases all phases $\theta(\omega_{k})$ by a small perturbation.
$F(\omega_{k})=\frac{1}{N}\sum^{N-1}_{t=0}{x_{t}e^{-j\omega_{k}t}}=A(\omega_{k})exp[j\theta(\omega_{k})]$ (10)
Decomposition
Empirical Mode Decomposition (EMD) [69] is used to produce a set of intrinsic mode functions (IMFs). This decomposition relies on extracting the energy linked to different time scales that are inherent to the data. EMD has been utilized as a method of data augmentation in [70] for the purpose of categorizing impact noise in vehicles. Another way to decompose time-series data is STL [71], which decomposes signals into seasonal, trend, and remainder:
$y_{t}=\tau_{t}+s_{t}+r_{t},\quad t=1,2,\cdots N$ (11)
Other works such as [72] and [68] focus on improving robustness with the help of STL components.
IV-B Structure-based Transformation
IV-B1 Shape Transformation
This type of data augmentation manipulates the spatial relationship between pixels. Image Cropping breaks the original image into pieces (Fig. 2d), and Geometric Transformation perturbs the structure within the original data (Fig. 2e).
Image Cropping
Image cropping selects an area from the original image and removes other parts of the image. It includes manual selection, random cropping, attention-based, aesthetics-based, and change-based approaches [73]. Attention-based approaches use certain attention scores to evaluate a subimage. Regions that contain the main topic or other types of information often have high attention scores. [74] uses the entropy, area ratio, and center distance of the image to evaluate the performance of the image segmentation and accordingly adjust the cropping parameters. [75] uses eye-tracking techniques to capture human fixations and thus infer the location of important contents on the image. Aesthetics-based approaches [76, 77, 78] are similar to quality evaluation. For example, [76] detects several subject regions and one background region, extracts features from the regions, estimates and combines posterior probabilities, and determines the quality score according to these probabilities. A2-RL [79] integrates image cropping into a sequential decision-making process and an aesthetics-aware reinforcement learning model to solve the best cropping problem. Change-based approaches such as [80] take into account the content that is removed and extract two sets of specially designed features to model what is changed after cropping.
Geometric Transformation
As one of the most classical image data augmentation methods, geometric transformation changes the spatial relationship of pixels within an image while keeping the value of each pixel untouched to create more data without extra sampling or labeling. Affine transformations like flipping, shearing, rotation, and translation manipulate the image while preserving the colinearity and ratios of distances. Image flipping reflects the image around its vertical axis, horizontal axis, or both axes, rotation turns the image by an angle about a fixed point. Translation means moving an image in a certain direction. It can introduce more variance in the position of target objects in the image. In non-affine transformation, the spatial relationship between pixels is further changed compared with affine transformations because the colinearity and ratio of distances are broken. Perspective transformation [81] mimics images taken from other views of the object. Image stretching [82] preserves or even enlarges the prominent contents in the image and compresses the background or unimportant foreground information. Distortion [34, 83] also stretches the pixels but often focuses on a certain region. PatchShuffle [84] divides an image into block matrices and shuffles the pixels within each block under a certain probability.
IV-B2 Syntax Transformation
We categorize syntax transformation as: Sentence Cropping that splits a sentence into many pieces, and Sentence Morphing that swaps the location of words or grammatical components. Some examples are presented in Table II.
Sentence Cropping
Sentence cropping [85] divides sentences into several small pieces. It often uses dependency trees to extract the relationships between subjects and objects of the sentence and produces multiple smaller sentences that focus on different parts of the original sentence. This operation breaks the syntactic solidity and semantic comprehensiveness of the entire sentence but preserves local syntactic tags and shallow semantic labels.
Sentence Morphing
These methods manipulate sentence structure by altering the position of words or terms. Random swap [11] randomly chooses two words in a sentence and swaps their position. This perturbs the structure and adds noise to the text data across paragraphs. [86] studies two types of syntactic augmentations: inversion and passivization. Inversion swaps the subject and object in a sentence, and the semantics might be different in most cases. Passivization converts a sentence to its passive form, and the meaning of the new sentence is typically consistent with the original one. Sentence rotation [85] borrows the idea of image rotation, where the image is rotated around its center. In sentence rotation, the root of the sentence is chosen as the center of the sentence, and other sentence fragments are connected to their center via the dependency tree.
TABLE II: Demonstration of some elementary operations for text data augmentation
Operation Sentence
Original I am an art student and I paint a lot of pictures. Many people pretend that they understand modern art.
Synonym Replacement [11] I am an art student and I draw a lot of pictures. Many people pretend that they understand contemporary art.
Similarity Replacement [43] I am an art student and I paint a lot of photos. Many people pretend that they understand modern literature.
Random Insertion [11] I am an art student and picture I paint a lot of pictures. Many appreciate people pretend that they understand modern art.
Punctuation Insertion [46] I am an, art student and: I paint a lot of pictures. Many people? pretend that they. understand modern art.
Random Swap [11] I am an of student and I paint a lot art pictures. Many people they that pretend understand modern art.
Random Deletion [11] I am an art student and I paint a lot of pictures. Many people pretend that they understand modern art.
Placeholder Replacement [47] I am an art _ and I paint a lot of pictures. Many _ pretend that they understand modern art.
Inversion [86] I am an art student and a lot of pictures paint me. Many people pretend that they understand modern art.
Passivization [86] I am an art student and I paint a lot of pictures. Many people pretend that modern art are understood by them.
IV-B3 Topology Transformation
Graph topology structure has immense information representation capacity. Naive Topology Perturbation methods modify a single node or edge component or consider a set of nodes as a whole and modify the substructure. Subgraphing extracts subgraphs from the complete graph. Graph Diffusion and Graph Rewiring are more sophisticated techniques that focus on adjusting the edges to obtain a better graph.
Naive Topology Perturbation
Due to their straightforward idea and easy implementation, node dropping and node addition are among the earliest augmentation approaches in graph learning. As Eq. 12 shows, Node dropping [49] randomly discards some nodes from the graph. Such an operation will also remove the attributes that these nodes carry and the edges on these nodes.
$\tilde{\mathcal{V}}=\text{Drop}(\mathcal{V}),\qquad\tilde{\mathbf{G}}=\{\tilde{\mathbf{A}},\tilde{\mathbf{X}}\}=\{\mathbf{A}[\tilde{\mathcal{V}},\tilde{\mathcal{V}}],\mathbf{X}[\tilde{\mathcal{V}},:]\}$ (12)
As for node addition, an example is MPNN [87]. It introduces a latent master node that is connected to all other nodes, which helps messages pass through the graph network. An edge describes the relationship between two nodes, and by perturbing these relationships through edge dropping and edge addition, we also introduce diversity to the structural information of a graph. DropEdge [88] perturbs the edge set by randomly setting a portion of the adjacency matrix to zero. It first produces a sparse matrix from a random subset of the original edges, then derives the augmented edge set by subtracting the sparse matrix from the original adjacency matrix. Edge masking in [51] constructs a Bernoulli distribution, and randomly masks edges under that probability. FairDrop [89] computes a mask according to the network’s homophily, and uses this mask to guide the dropping process to remove unfair connections. AD-GCL [90] applies a GNN-augmenter to learn how to drop edges so that the mutual information of the augmented graph and that of the original graph is minimized.
Subgraphing
Subgraph cropping is based on the idea that graph semantics are partially represented in local structures [49]. Some subgraphing methods, such as those used in GCC [91] and SUGAR [92] use random walk to sample a subgraph from the original graph. GrpahCrop [93] applies a node-centric strategy to crop connectivity-preserved subgraphs and expand the dataset with these augmented subgraphs. These subgraphs are able to maintain some topology characteristics of the original graph. Similarly, SUBG-CON [94] uses the Personalized PageRank algorithm to select the important nodes. NeuralSparse [95] proposes to obtain the subgraph by learning. It uses a sparsification network to generate sparsified subgraphs. Subgraph Substitution perturbs the substructure as a whole of a graph, which is especially beneficial for molecular graphs. A molecule graph represents atoms with nodes and chemical bonds with edges. Functional groups, which often decide the chemical characteristics of molecules, can be represented by substructures. So, by wisely sampling substructures, the global semantics of a molecule graph can be well contained. MoCL [96], for example, utilizes substructure information as one of its augmentation schemes. It performs graph augmentation by replacing functional groups, effectively introducing reasonable noise. GEAR [97] also employs such a technique. It augments molecule graphs by environment replacement, introducing noise from environment subgraphs.
Graph Diffusion
Traditional message passing in graph neural networks only leverages one-hop connections, while the edges of a real graph are often noisy or arbitrarily defined, and many relationships in the networks are indirectly contained in higher-order neighborhoods. To address this issue, GDC [98], as in Eq. $\mathbf{\tilde{A}}=\sum_{k=0}^{\infty}\theta_{k}\mathbf{T}^{k}$ , proposes to calculate a series of transition matrices $\mathbf{T}^{k}$ by means such as random walk and combines these transition matrices with weights $\theta_{k}$ to obtain a new adjacency matrix $\mathbf{\tilde{A}}$ and thus generate a new graph that directly contains some node connection hidden in the original graph. The new graph can also be understood as being processed by a denoising filter because graph diffusion smooths out the neighborhood over the graph [98]. This technique is widely used by self-supervised approaches to transform the graph and generate different views for contrastive learning, such as MVGRL [99], SelfGNN [100], and MERIT [101].
Graph Rewiring
Graph rewiring, as presented in Eq. 13, is initially proposed as randomly swapping edges [102] to improve a network’s robustness and mitigate malicious attacks. Some later research such as smart rewiring [103] takes the degrees of chosen nodes when swapping their edges. Stochastic Discrete Ricci Flow [104] is a curvature-based graph rewiring method, which in each step adds an edge that can improve the most negatively curved edge, and then removes the edge with the most positive curvature. Its results demonstrate the ability of graph rewiring to alleviate the over-squashing problem. DiffWire [105] introduces the CT-LAYER and GAP-LAYER in a GNN model, which respectively learns the commute times and spectral gap of the graph. It rewires the adjacency graph so that the graph’s spectral gap is minimized, and by doing so tries to predict the optimal graph structure for the downstream task. DHGR [106] first learns a similarity matrix from the original graph. Then it rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones. Such an operation can effectively improve the homophily of the graph and thus improve GNN performance. Following DHGR [106], HDHGR [107] computes multiple similarity matrices under different meta-paths and thus generates several rewired meta-path subgraphs.
$\vspace{-2mm}\mathbf{\tilde{A}}=\overbrace{\mathbf{A}\odot(\mathbf{1}-\mathbbm{1}_{r})}^{\text{edge removal}}+\overbrace{(\mathbf{1}-\mathbf{A})\odot\mathbbm{1}_{r}}^{\text{edge addition}}$ (13)
IV-B4 Table Structure Transformation
Table Subsetting
Inspired by image cropping [73, 80], SubTab [108] divides a table into multiple subsets by splitting its columns. Each subset contains only a percentage of all features. The model is forced to learn the representation of the samples by these features and is required to reconstruct all the features from the subset.
IV-B5 Sequence Transformation
Unlike Value Perturbation, sequence perturbation is implemented on the X-axis, either on the time domain or the frequency domain. Window Slicing splits the time-series data into pieces. It is also the basis of many Sequence Morphing methods because they apply varied operations on these slices.
Window Slicing
Inspired by image cropping, which is widely used for image data augmentation, MCNN [109] proposes window slicing. It extracts slices from the entire sequence as new training samples. Each slice is assigned with the label of the original sequence.
Sequence Morphing
Based on Window Slicing, [110] extends or compresses the generated spans, which is called window wrapping. This is equivalent to speeding up or slowing down a span of the sequence. Vowel stretching [111] is a special kind of window wrapping technique. It only perturbs the sequence by prolonging the vowels. This is meant to be consistent with how children speak. [67] propose to randomly permeate these slices and concatenate them as a long sequence. Similar processing is also called Random Shuffling and is used in [112]. VTLP [113] applies window wrapping in the frequency domain. It adds a small distortion to the central frequency. SpecAugment [114] also proposes to augment the frequency domain. It generates masks of length sampled from a uniform distribution and uses each of them to mask consecutive frequency channels.
IV-C Value-structure Transformation
IV-C1 Image Value-structure Transformation
Policy-based methods search for the optimized augmentation policy that includes both value-based transformations and structure-based transformations, as demonstrated in Fig. 2f. Prompt-based Image Editing employs a text-to-image model to change the content, background, or style of the image as humans desire. When implementing the augmentation, both coloring and the positions of pixels are changed (Fig. 2g).
Policy-based Data Augmentation
AutoAugment [13] represents data augmentation as a decision-making process and uses reinforcement learning to search for the best operations. In AutoAugment, a policy includes a group of basic data augmentation operations with probabilities and magnitudes. AutoAugment processes images under the policy, evaluates these augmented images on downstream networks, and rewards a policy generator for better performance. Fast AutoAugment [115] proposes a more efficient searching algorithm for augmentation policies, which works significantly faster than AutoAugment and thus allows a larger number of sub-policies, which should benefit the generalization performance. RandAugment [116] removes the need for a separate search phase on a proxy task, and makes the training process less complicated and less expensive. Population Based Augmentation [117] learns an augmentation schedule rather than a fixed policy as in AutoAugment to improve efficiency. These methods combine the benefits of both value-based transformations and structure-based transformations and further discover the potential of data augmentation with individual image data.
Prompt-based Image Editing
Visual generative models have been rapidly evolving in recent years. These models are typically conditioned on text embeddings, allowing flexible image editing by modifying textual descriptions of images and prompting the models. DA-Fusion [118] learns fine-grained concepts, replaces the original category labels, and uses the Stable Diffusion model to generate new images. SaSPA [119] uses GPT-4 to generate prompts based on the sample’s meta-class, then inputs the prompts, reference images, and detected edges to the Diffusion model to generate augmented images.
IV-C2 Text Value-structure Manipulation
Hierarchical Data Augmentation explicitly augments both value and structure information of text data. Back-translation Augmentation and Paraphrase Augmentation use neural networks to generate sentences that are semantic-preserving but different from original sentences in both wording and syntax.
Hierarchical Data Augmentation
Hierarchical Data Augmentation (HDA) [120] augments text data at both word and sentence levels. It evaluates the roles of words and sentences by training a hierarchical attention network to obtain the attention values of the text data at both levels. Attention values help decide the importance of components and augment the data by cropping and concatenating these components.
Back-translation Augmentation
Back-translation Augmentation. Back-translation [121] means translating text data from one language to another, or a sequence of other languages, and finally translating back to the original language. Semantic information is preserved, but the translation process can introduce some variance in syntax and wording. Such transformation can help the model learn more flexible expressions that convey the same meaning.
Paraphrase Augmentation
AugGPT [122], for example, prompts the state-of-the-art ChatGPT model to rewrite sentences in the dataset and preserves dataset coherence after data augmentation. Taboo [123] requires paraphrasing without using some frequently used words, Chain [124] paraphrase one sentence multiple times, and Hint [124] provide some examples, to encourage the LLM response otherwise. DoAug [125] constructs a preference dataset based on the diversity of paraphrased sentences and trains the LLM with the DPO to boost its generation diversity.
V Multi-instance Level Data Augmentation
Multi-instance augmentation combines value and structure information from different samples. Specific operation varies depending on the data modality in which it is applied.
Multi-instance level image augmentation combines the visual information, i.e., pixels, from different images. It either mixes the colors of images or puts their patches together. Multi-instance level text augmentation integrates the linguistic information from different sentences. It either mixes the sentence at the word level or combines fragments from multiple sentences. Multi-instance level graph augmentation mixes the network information of different graph components or graph instances. It either mixes the attributes of different nodes or combines both the topology and attributes of graphs. Multi-instance level tabular augmentation combines the information of different lines or tables. Most methods mix the value representations across different lines, while it is also possible to combine two heterogeneous tables. Multi-instance time-series augmentation mixes values from different sequences, with few methods combining sequences using structural relationships.
V-A Value-based Mixture
V-A1 Pixel Mixture
Image Mixup computes an arithmetical mixture of the color from different images and Neural Blending resorts to neural networks to enhance such process.
Image Mixup
Mixup [126] mixes two random images on the input layer by calculating their interpolation pixel-wise. Pixels from two original images each contribute a portion to the new image:
$\vspace{-2mm}\mathbf{\tilde{x}}=\lambda\mathbf{x}_{i}+(1-\lambda)\mathbf{x}_{j},\qquad\mathbf{\tilde{y}}=\lambda\mathbf{y}_{i}+(1-\lambda)\mathbf{y}_{j}$ (14)
where $(\mathbf{x}_{i},\mathbf{y}_{i})$ and $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two labeled images from the dataset, and $\lambda\in[0,1]$. As a variant, manifold mixup [127] is operated on the hidden layers of the neural network. Most mixup methods combine two different images, while AugMix [128] applies three operations to one image, and combines the three augmented images to generate a new sample. In some approaches, the mixing operation varies in strength across regions. For example, SmoothMix [129] applies a soft-edged mask when mixing. Most mixup methods create new samples with soft labels, as in Eq.14, while there are some exceptions. Sample Paring [130], for example, uses the label of one of the original images, and that of the other image is left unused. ReMix [131] addresses the class imbalance problem by assigning the minority class label to augmented data.
Neural Blending
AutoMix [132] integrates a Mix Block in the training process. It builds a bridge between the selection of mixup policy and the optimization of the model. In AutoMix, the augmentation module is optimized along with the classification feature map, and the mixup functions and mixing ratio will updated according to the classification loss. Another image blending technique is proposed as Smart Augmentation [133]. At the input end of Smart Augmentation, two random samples are fed into the network through two channels, and a CNN sub-network yields a one-channel output with the same shape as the input, which means one new image is generated based on the two input images. The augmented image is then used for downstream tasks, and the augmentation sub-network adjusts to the task performance accordingly.
V-A2 Text Representation Mixture
Unlike image data, whose coloring information is expressed by numerical values and can be interpolated at the input layer, words are difficult to interpolate directly, so text interpolation is usually done at the embedding layer.
Text Mixup
The adaptation of Mixup in sentence classification is studied in [134], which proposes wordMixup and senMixup. WordMixup interpolates a pair of word embeddings and then feeds the interpolation to the sentence encoder. SenMixup, on the other hand, interpolates a pair of sentence embeddings. The embedding vector produced by WordMixup or SenMixup is then passed to the softmax layer to produce the target distribution of the classification task. The use of mixup in text data augmentation has also been extensively studied in many research, such as nonlinear Mixup [135] and DoubleMix [136]. GPT3Mix [137] makes use of the latest large language model to mix sentences. By inserting two source sentences in a prompt template that describes the instruction to the model, we can get a mixed sentence of high quality with a soft label.
V-A3 Node Mixture
Feature or label information from different nodes can be mixed by Graph Propagation, which is often used for tasks such as node classification.
Graph Propagation
Propagation refers to spreading information from one node to other nodes along the graph. These methods include feature propagation and label propagation. Feature propagation shares node features with other nodes. After feature vectors of some nodes are masked out, GRAND [50] further augments the node features by randomly propagating them through edges. MV-GCN [138] first creates three different views from the original graph, propagates the features in each, and then combines the three representations. Label propagation algorithm (LPA) [139] propagates labels from labeled nodes to all unlabeled nodes according to their proximity through areas with dense unlabeled nodes. AutoGRL [140] applies LPA in their semi-supervised framework to label more nodes utilizing a few ground truth nodes. [141] performs two types of label propagation, correcting the base predictions and smoothing the final prediction, respectively. Though not in a typical graph-based scenario, [142] constructs a nearest neighbor graph on the classifier’s manifolds. The affinity matrix, analogous to the adjacency matrix in this case, is constructed based on the $k$ nearest neighbors of each node’s embeddings, and this matrix decides how labels propagate.
V-A4 Cell Value Mixture
Due to its well-aligned structure, tabular data is very suitable for interpolation.
Tabular Mixup
SMOTE [37] is among the first works to study data interpolation, even before the proposal of Mixup [126]. It addresses the problem of class imbalance by oversampling from the minority class. The samples are selected within the $k$ nearest neighbors in the feature space and new synthetic data are produced by a random interpolation operation. Since tabular features are often nominal, two variants, namely SMOTE-NC and SMOTE-N, are also proposed along with the initial version of SMOTE [37]. As argued in [143], SMOTE is subject to unrealistic samples when interpolating in a sparse feature space. To solve this problem, TAEI [143] first encodes the sparse data into a dense latent space, interpolates the new data, and then decodes them to obtain the original data plus the augmented data. Contrastive Mixup [144] applies mixup for tabular data in a self- and semi-supervised learning framework. SAINT [145] combines CutMix [146] with contrastive learning to learn from tabular data. ExcelFormer [147] proposes two schemes, namely Feat-Mix and Hidden-Mix. They interpolate data in the input layer and the hidden layer, respectively.
V-A5 Sequence Value Mixing
Methods that combine information from different sequences are referred to as sequence averaging or sequence mixing. DTW Barycenter Averaging (DBA) [148] revolves around the averaging problem for Dynamic Time Warping. It develops a global technique for computing the average of a set of sequences and avoids using iterative pairwise averaging. In Weighted-DBA [149], instead of each time series contributing equally to the final average, some can contribute more than others, so that an infinite number of new examples from any set of given time series can be generated. [150] applies both interpolation and extrapolation in the feature space to augment time-series data. It first uses a sequence autoencoder to learn a feature space from unlabeled data. The data is then encoded and augmented with additive noises, interpolation, and extrapolation. Finally, the data in feature spaces is decoded and classified. Extending sequence mixing to the frequency domain, [151] applies the Equalized Mixture Data Augmentation (EMDA) to create new data. It computes the weighted average of two randomly chosen spectrograms with the same label.
V-B Structure-based Combination
V-B1 Image Structural Combination
Image Patching puts together small fragments of many images or puts a part of an image over a second image.
Image Patching
RICAP [152] randomly crops four images, and patches them to construct a new image. CutMix [146] makes use of two images. It is developed based on Cutout [9], but the removed region is replaced by a patch from another image. Attentive CutMix [153] further enhances the CutMix method by adding a feature extraction module that selects an important or representative region, and pastes this attentive patch onto the other image. CowMask [154] introduces patching with irregular shape rather than rectangular.
V-B2 Text Structural Combination
Text Fragment Merging combines text data by merging text fragments from one sentence into another.
Text Fragment Merging
GECA [155] aims to remove a fragment from a sentence and populate it with another fragment that occurs in some common environment. Subtree Swapping [156], Substructure Substitution [157], and TreeMix [158] use tree structures to analyze the part of speech or role in the sentences and thus make reasonable swaps. SSMix [159] replaces a low saliency segment in a sentence with a segment with high saliency from another sentence and mixes sentence labels according to the ratio of their span length in the combined sentence.
V-B3 Table Concatenation
Contrary to single-instance level tabular augmentation, which lacks structural information to use, multi-instance level tabular augmentation suffers from strict and fixed table structures. Columns differ in their nomenclatures and representations, so it is hard to combine and learn from multiple distinct tables. If these restrictions are relaxed, models can learn across different tables, be more transferable, and better predict unseen tables.
Structure Relaxation
TransTab [160] suggests that multiple tables could share partially overlapped columns in the real world, but traditional methods often fail to fully utilize the data due to the removal of non-overlapping columns and mismatched samples when cleaning the data. TransTab first divides the table into three parts, each containing categorical, binary, and numerical data. These three types of data undergo different operations, are all converted to embeddings, and are again integrated into a new and encoded table. The processed tables can be joined in many ways, depending on the application scenarios. The augmented table contains information from many original tables and helps the model learn across them.
V-C Value-structure Mixture
V-C1 Image Value-structure Mixture
Diffusion Models take textual prompts as inputs and can perform a flexible and smart mixture of existing images.
Prompt-based Mixing
Conditioned on textual inputs, Diffusion Models can generate combined images by using textual descriptions of multiple images. Diff-Mix [161] uses text embeddings of different categories to replace the foreground to align with the target classes. Diff-II [162] calculates the inversion noise of two images, conducts random circle interpolation, and then denoises with different prompts.
V-C2 Graph Value-structure Mixutre
Graph Mixup refers to mixing graph nodes or graph instances. This often involves both node attribute values and the topology structures.
Graph Mixup
Due to the irregular connectivity and various topological structures of graph data, it is hard to directly combine nodes or graphs. For node-level mixup, one solution to this problem is temporarily ignoring the edges and mixing node features in latent space. GraphMix [163] interpolates the hidden states and labels of input nodes. A topology-involved node mixup method is proposed in [164]. It randomly pairs nodes and includes the nodes’ local topology representations when interpolating them. S-Mixup (Structural Mixup) [165] considers neighboring edges when pairing nodes but works in the input space and discusses how to connect the node with the original graph with edges. It matches nodes based on GNN classification confidence and connects with high-gradient edges to the mixed node. For graph-level mixup, ifMixup [166] adds dummy nodes so that two graphs have the same number of nodes. These dummy nodes have all-zero feature vectors and are not connected to any other nodes, or in other words, the weights of edges on dummy nodes are zeros. S-Mixup (Graph Mixup with Soft Alignments) [167] also aligns graphs first before mixing them. It obtains an assignment matrix by a soft assignment and uses the assignment matrix to transform a graph so that every node corresponds with a node in the other graph. Compared with the above approaches, G-Mixup [168] proposes a more sophisticated way to interpolate graph data. It first uses graphs within the same class to estimate a graphon, i.e. graph generator:
$\vspace{-2mm}\mathbf{\mathcal{G}}\rightarrow W_{\mathbf{\mathcal{G}}},\mathbf{\mathcal{H}}\rightarrow W_{\mathbf{\mathcal{H}}}$ (15)
The graphon $W:[0,1]^{2}\rightarrow[0,1]$ is a matrix, where $W(i,j)$ describes the probability of an edge between nodes $i$ and $j$, and graphs of the same class are considered to be generated from the same graphon. Graphons of two classes of graphs are then interpolated to produce a mixed graphon:
$\vspace{-2mm}W_{\mathbf{\mathcal{I}}}=\lambda W_{\mathbf{\mathcal{G}}}+(1-\lambda)W_{\mathbf{\mathcal{H}}}$ (16)
By sampling from the graphon mixup $W_{\mathbf{\mathcal{I}}}$, we acquire a series of graphs with the same soft labels:
$\vspace{-2mm}\mathbf{y}_{\mathbf{\mathcal{I}}}=\lambda\mathbf{y}_{\mathbf{\mathcal{G}}}+(1-\lambda)\mathbf{y}_{\mathbf{\mathcal{H}}}$ (17)
During graphon estimation, G-Mixup aligns the original node features and obtains the graphon node features average pooling, which are used as node features of the generated graphs. DAGAD [169] mixes graph representations by performing random permutations, demonstrating that data augmentation is helpful in scenarios such as graph anomaly detection.
VI Dataset Level Data Augmentation
Dataset-level data augmentation methods learn the overall data distribution and often generate new data that is consistent but diverse in terms of both value and structure information.
VI-A Vanilla Generation
A straightforward way to produce augmented data from the entire dataset is to learn the data distribution of existing datasets with neural networks and generate new samples that fall in that distribution. Some typical solutions are Autoencoder (AE), Generative Adversarial Networks (GAN), Large Language Models (LLMs), and Generative Diffusion Models.
VI-A1 Autoencoder
An Autoencoder (AE) learns from the dataset and extracts the essential structure, which can be used to generate a more robust dataset. It is widely used for tabular data augmentation. [170] tries out four types of autoencoder, concluding that all of these autoencoders can generate data sets that produce better task performance, while the variational autoencoder (VAE) provides the most robust result. Some other works [171, 172] also use VAE as their data generation approach. SDAT [171] adds extra noise to the latent space of its VAE to obtain augmented samples. These samples, along with the original ones, are then trained under a semi-supervised framework. [172] combines VAE with active learning. The samples are iteratively selected by active learning, labeled by an oracle, reconstructed by the VAE, and discriminated and classified.
VI-A2 Generative Adversarial Networks
The rise of Generative Adversarial Networks (GANs) [173] makes it possible for humans to generate new images on a large scale using computers. This nature also leads to a new path of data augmentation: creating new data samples instead of modifying or combining old ones. It can produce considerably more augmented data and is especially useful for oversampling and alleviating the problem of class imbalance.
For image data, GANs have now been widely used as a data augmentation technique [12], especially in the medical field [174, 175], where data is to some extent scarce. Apart from simply enlarging the dataset, GAN can be used in other ways such as removing background noise and cleaning the dataset [176, 177, 178]. It can also generate different views or poses of the object based on the original image [179], and thus add more variation to the dataset.
Tabular data augmentation is another case where GANs play an important role. Table-GAN [180] mainly considers the privacy issue, and experiments show that the augmented table given by GAN is compatible with downstream machine learning models and produces similar performance. CTGAN [181] focuses on the class imbalance problem and proposes a conditional generator that accounts for the imbalance in categorical columns. [182] shares a similar goal and method with CTGAN [181] but employs a more sophisticated loss function that better incorporates the discriminator and the auxiliary classifier. ITS-GAN [183] proposes a framework that aims to preserve functional dependencies in tables. It models functional dependencies with autoencoders and includes a term to characterize the set of functional dependency constraints in the generator loss function. Arguing that GANs are difficult to train on small datasets, ConvGeN [184] designed a new mechanism that requires the GAN’s generator and discriminator to cooperate, that the generator generates batches of a convex combination of minority neighborhood samples and the discriminator classifies these augmented minority samples against majority batches.
GANs are also widely used to generate virtual time-series data. [185] uses an RGAN and an RCGAN to produce realistic medical time-series data. TimeGAN [186] combines supervised and unsupervised losses to train its autoencoding components and adversarial components. [187] propose WaveGAN and SpecGAN to generate time-series data. WaveGAN is used on the waveform, while SpecGAN is used on the spectrogram. SpecGAN demonstrates how GAN is applied to generate time-series data on the frequency domain.
VI-A3 Large Language Models
Recent advances in large language models (LLM) and generative pre-training (GPT) [188] models have demonstrated unprecedented power in text understanding, classification, and generation tasks. Apart from supporting intelligent chat-bots that help people in their daily lives, the capacity of these state-of-the-art language generation models establishes a new paradigm of text data augmentation.
LAMBADA [189] is among the first attempts to use a state-of-the-art language generator to synthesize new text data for data augmentation, and its procedure is as follows. LAMBADA first trains a baseline classifier from a small dataset, and then it fine-tunes the GPT-2 [190] model. A set of labels is thereafter fed to the fine-tuned language model, which generates a large dataset of labeled sentences. These sentences are evaluated by the classifier trained prior according to its confidence score, and top-ranked sentences are retained to compose the synthesized dataset. Similarly, [191] integrates GPT-2 in its text data augmentation process and improves the performance of active learning on tiny datasets. It uses the language model to generate a tree of tokens as candidates and construct a sentence from them. The set of new sentences is repeatedly refined by active learning, and the outcome is an augmented text dataset.
Some other research also applies GPT-2 to effectively augment low-resource text datasets, such as [192] and [193]. Despite their success, both of them point out the need for more sophisticated methods to generate text data. [192] for example, argues that human-in-the-loop for relabeling might be rewarding. [193] adopts human-in-the-loop for seed selection and further discusses the potential benefits of injecting expert knowledge for specialized domains. Notably, Data Boost [194] proposes a reinforcement learning framework to guide the process of text data augmentation with GPT-2 and succeeds in generating data of high quality and high similarity to the original data.
The rise of GPT-3 [195] and later LLMs is now providing more possibilities for text generative augmentation. Their use and improvements are studied in the latest research such as [196] and [197].
VI-A4 Generative Diffusion Models
Despite that AEs and GANs can all be used to generate graph data, they face many problems in practice. VAE models require things like a massive graph-matching process, and GANs on graphs easily fall into mode collapse [198]. So, graph data augmentation resorts to another recently proposed technique, namely, Generative Diffusion Models.
The denoising diffusion model [199] is a generative paradigm that has become popular in recent years. Though not yet widely applied as an approach to graph data augmentation, the Generative Diffusion Model boasts enormous potential in graph generation and can possibly be adopted as a promising data augmentation technique. To implement diffusion models on graph data, some properties of the graph data must be properly treated. For example, diffusion models often rely on a continuous Gaussian noise process, but the graph structures are discrete [200]. DiGress [201] designs a discrete denoising diffusion model for graph generation and achieves state-of-the-art performance on molecular and non-molecular datasets. The use of Generative Diffusion Models for data augmentation especially benefits studies on molecules, proteins, and materials, whose data are often in the form of graph instances.
VI-B Exogenous Generation
Apart from the source dataset, we can also refer to exogenous information for data augmentation. Exogenous information is perceived by introducing external datasets or using knowledge learned by human experts.
VI-B1 Introducing Exogenous Datasets
Neural Style Transfer
Image datasets vary in artistic styles, and neural style transfer [202] introduces style information from an external dataset when augmenting an entire dataset. It learns the artistic style from a group of image data and uses this artistic style to depict the topic of images from another dataset. As is [202], neural style transfer can be used to mimic the drawing styles of particular artists. Such an augmentation process can help the model learn a more general representation of the topics. [203] proposes a more realistic use of neural style transfer. It transfers the weather or illumination from reference images to the target dataset. This is particularly useful because it is hard to collect sufficient images under certain weather.
Table Retrieval
Tabular data populates the web and by retrieving consistent and related tables, it is possible to generate an augmented table that incorporates extended information. A key consideration of Table Retrieval Augmentation is to design appropriate ways to search for ideal tables. Santos [204] leverages external knowledge bases, YAGO 4, and focuses on designing a union search technique to match desired target tables. HYTREL [205] predicts table similarities by building hypergraphs on top of the tables and uses a hypergraph encoder to calculate the encoding. LEKA [206] uses an LLM to extract the table’s descriptive information, retrieve it in the Kaggle database, and harmonize tables from different sources.
VI-B2 Leveraging Expert Knowledge
Here we introduce three examples of how humans utilize expert knowledge for image, tabular, and time-series data augmentation
Computer Graphic Modeling is to leverage expert knowledge on how objects in the existing dataset look like and the power of computer software to generate new image samples [207, 208, 209]. These datasets have several advantages over ordinary image datasets. First, it requires little human effort to collect data. Second, by applying different aspects, poses, and lighting conditions, computer graphics modeling can generate images with high diversity. Third, these images are all fully labeled with accurate ground truth.
Relational Structure Construction can enhance the potential of tabular data, which is limited by its simple nature compared with data from other modalities. Due to the lack of spatial or relational relationships, many popular augmentation approaches and network configurations are not suitable for tabular datasets. A solution is to construct and introduce a relational structure based on the tabular dataset. GOGGLE [210] proposes a solution to this problem. It learns the relational structure and corresponding functional relationships from tabular datasets and uses this basis to generate new samples. In other words, it learns a graph from the features in the table and obtains augmented data from that graph. PET [211] depicts multiple tabular data instances as a hyper-graph and uses the labels to help construct hyper-edges. These methods are based on external assumptions and introduce extra structural information beyond the tabular dataset.
Statistical Generation is a technique that predicts values in the future. Some methods view time-series data outside the existing data and from a statistical perspective. They are based on human knowledge of statistics and can predict how the value changes in future timestamps. [212] uses a statistical algorithm called LGT (Local and Global Trend) to forecast paths of the time-series data and samples from them to get extra data. GRATIS [213] employs mixture auto-regressive (MAR) models to generate collections of time series and examine the variety and extent of the produced time series within a feature space for time series.
VII Discussions
Up to now this survey focuses on wherefrom the information in the augmented samples is derived. And there are many other perspectives to discuss data augmentation and understand how it works.
VII-A Augmentation Evaluation Metrics
Most research on data augmentation uses the models’ performance on downstream tasks to evaluate the effect of data augmentation [13]. However, this criterion depends on many extraneous factors and may fail to reflect the influence of data augmentation methods on the data. So, there is a need for a set of evaluation metrics that can assess specific considerations of data augmentation. Two of the most critical metrics are affinity and diversity [161, 162, 214, 215, 216, 217, 125].
Affinity is also referred to as consistency, correctness, faithfulness, quality, similarity, or validity. It reflects if the augmented sample is semantically consistent with the original sample and preserves the original label. Model-centric evaluation [215] defines affinity as the ratio of model $m$’s performance on augmented validation dataset $D^{\prime}_{val}$ and that on the original validation dataset $D_{val}$, given by
[…]
where $\text{ACC}(m,D)$ stands for model $m$’s performance on evaluation dataset $D$. Data-centric evaluation, on the other hand, defines affinity as how much the augmented dataset deviates from the original dataset in the feature space. For example, [214] proposes Fréchet BERT Distance (FBD) to evaluate the deviation of two textual datasets, defined as:
[…]
where $m$ is the mean vector and $C$ is the covariance matrix.
Diversity reflects the breadth of patterns in the dataset and exposing the model to a diverse training dataset can improve its robustness. [218] considers a dataset as an ellipsoid and defines its diversity as the isocontour radius of the ellipsoid:
[…]
where $H$ is the dimension of the feature vectors and $\sigma_{i}$ is the standard deviation along the $i$-th feature. [219] proposes to evaluate dataset diversity by calculating the dispersion:
[…]
where $d(x_{i},x_{j})=1-\frac{f(x_{i}){\cdot}f(x_{j})}{\|f(x_{i})\|\|f(x_{j})\|}$.
Some empirical studies including [219], [217], and [215] investigate how affinity and/or diversity influence the effectiveness of data augmentation, and their real-world experimental results demonstrate the model performance takes the most advantage of data augmentation that achieves both high affinity and diversity.
VII-B Selecting the Target
VII-B1 On Selecting Pixels
For Pixel Erasing, selecting target pixels is a necessity because it is impossible to mask out all pixels in the image. If erasing is performed by randomly selecting an area in the image, it is called Image-aware Random Erasing (IRE) [10]. In some tasks, the location of the target in the image is known, so, random erasing can be restricted in the object’s bounding box, known as Object-aware random erasing (ORE) [10]. Further argued in [220], the erased area within the bounding box should be limited to a manageable size and a part of the less informative background should also be erased. [220] proposed Region-aware Random Erasing (RRE), which introduces ORE with upper bounds for width and height in the bounding boxes, and then randomly erases the background. The same consideration occurs when performing Image Cropping and Mixup, and some importance-based strategies are used to preserve important contents in the images. For Image Cropping (IC), [74] uses entropy as the importance score to select the most importance regions, and [75] uses eye-tracking technique to locate important content. For Image Mixup, Co-mixup [221] and Puzzle mix [222] both use saliency to decide the mixing strength of different regions, potentially leading to better mixing performances. A summary of how these methods select target pixels are presented in Table III.
TABLE III: Region selection strategy for Pixel Erasing, Image Cropping and Mixing methods.
Name Ref Augmentation Strategy
Random Location-based Importance-based
Cutout [9]
Image-aware RE [10]
Object-aware RE [10]
Region-aware RE [220]
Hide-and-seek [38]
Grid-Mask [39]
Entropy-bases IC [74]
Gaze-based IC [75]
Co-mixup [221]
Puzzle mix [222]
VII-B2 On Selecting Words
Most value-based text transformation methods randomly select the word to be replaced, added, and deleted, but some methods considers the role a word plays in the sentence when deciding whether to impose perturbation.
For example, TEXTFOOLER [223] ranks words by their importance and replaces the important words with their synonyms. Syntax-aware data augmentation [48] argues that most previous text data augmentation approaches are only computation-motivated and cannot assure the good quality of the generated text. Instead of simply modifying words with the same probability, Syntax-aware data augmentation calculates the probability of being modified for each word based on its syntactic role in the sentence. Specifically, it utilizes a dependency tree to calculate the importance of each word, i.e. the depth in the tree. A word far from the root is considered less important syntactically and is more likely to be removed (by means of dropout and blanking) or replaced during the augmentation process. Instead of using a syntax tree to assess word importance, Selective Text Augmentation (STA) [224] focuses more on the statistical and semantic features of words to describe the role of words. It proposes two perspectives to view words in text classification tasks, namely, statistical correlation, which means how frequently a word co-occurs with a certain category, and semantic similarity, which measures the semantic overlap of the given words and the category label. Using the metrics as axes, STA divides the roles of words into four quadrants: gold, venture, bonus, and trivial. Words with different roles are treated differently when imposing augmentation operations on the sentence. For example, gold words will be kept in selective deletion and synonyms of gold words will be added to the sentences in selective insertion. A summary of how these methods select target words is presented in Table IV.
TABLE IV: Word selection strategy of text data selection methods.
Name Ref Augmentation Strategy
Random Role-based Learned
EDA [11]
- [43]
TEXTFOOLER [223]
LetterTran [225]
AEDA [46]
INSNET [226]
LeCA [227]
GBS [228]
- [47]
STA [224]
Subtree Swapping [156]
Substructure Substitution [157]
TreeMix [158]
SSMix [159]
VII-B3 On Selecting Nodes and Edges
For graph data augmentation, a key problem is selecting target nodes and edges to operate on. Node centrality acts as an important factor that is both meaningful and convenient, supporting selection strategies of many graph data augmentation methods. GCA [229] uses node centrality to calculate the masking vectors and rewiring probabilities and thus decide which node to mask and which edge to rewire. GrpahCrop [93] applies a node-centric strategy to preserve connectivity in subgraphs. SUBG-CON [94] uses the Personalized PageRank centrality to select the important nodes for subgraphing. Smart rewiring [103] takes the degrees of chosen nodes when swapping their edges. Another useful factor is node similarity. DHGR [106] learns a similarity matrix from the original graph and rewires the graph by adding edges between high-similarity node pairs and removing edges from low-similarity ones. Similarly, RobustECD [230] and M-Evolve [231] also consider node similarity for edge rewiring.
TABLE V: Node and edge selection strategy for grpah data augmentation methods.
Name Ref Augmentation Strategy
Random Centrality Similarity
Attribute masking [49]
GRAND [50]
BGRL [51]
GCA [229]
GrpahCrop [93]
SUBG-CON [94]
Smart rewiring [103]
DHGR [106]
RobustECD [230]
M-Evolve [231]
VII-C Leveraging Augmented Samples
A basic use of data augmentation is to enlarge the dataset, especially in data-scarce scenarios, where the augmented samples are appended to the training dataset. Beyond this paradigm, data augmentation provides many other possibilities to support deep learning. A frequent usage of augmented samples is to construct different views for contrastive learning.
For example, Rock-ViT [232] performs rotation, flipping, blurring, contrast enhancement, irregular noise, and microscope dark-field noise on sandstone microscopic images. It optimizes the vision transformer with both cross-entropy loss and contrastive loss. When calculating the contrastive loss, Rock-ViT uses two random augmentations to build positive pairs. Instead of simply calculating contrastive loss between positive and negative sample pairs, DualCL [233] augments text samples by including the label in the sentence. Specifically, for an original sentence “[CLS] love this movie” in a sentiment classification dataset, DualCL converts it to “[CLS] POS NEG love this movie”, where “POS” and “NEG” denote the positive and negative labels. Then it repels the embeddings of the “[CLS]” token and the “NEG” token, since this sample is possible and thus “[CLS] should also be positive. Conversely, it attracts the embeddings of the “[CLS]” and “POS” tokens. Contrastive learning is a common practice in graph deep learning, and graph data augmentation plays a critical role in many graph deep learning pipelines. GraphCL[49] uses Node Dropping and Edge Perturbation to build different views, and MVGRL [99] applies graph diffusion. Different views of an original graph are used as positive pairs and a contrastive loss function is used to enforce maximizing their consistency.
VIME [56] expend the usage of data augmentation beyond contrastive learning. It incorporates tabular data augmentation in both self-supervised and semi-supervised learning frameworks. For self-supervised learning, it corrupts the table with a mask and optimizes the model with table reconstruction loss and mask estimation loss. For semi-supervised learning, it calculates the consistency loss between masked views.
VII-D Harnessing Latest Techniques
Early data augmentation methods rely on hand-craft manipulations, which are constrained by effectiveness and efficiency. AutoAugment [13] begins to introduce reinforcement learning to find optimal augmentation policies, enabling automatic search for desired augmentation strategies. Generative models, such as LLMs and diffusion models, possess outstanding capacities to capture sample semantics and produce new samples with high flexibility and high quality [234, 162, 122]. However, data augmentation is sometimes only a preprocessing step in a deep learning framework, and introducing generative models might introduce extra computation overhead. How to balance the tradeoff between the reward and cost of using these advanced techniques remains an open but critical problem.
VIII Conclusion
This survey presents a comprehensive summary of data augmentation techniques across five data modalities and proposes a modality-independent taxonomy from a data-centric perspective, which focuses on from where the augmented data is derived. It further enumerates related research papers on these data augmentation methods and annotates them with descriptive information in detail. Finally, it discusses several critical issues related to data augmentation that can help us better understand this technique.